Circular No. 8878

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 T2 (KOWALSKI) R. A. Kowalski reports his discovery of a comet on Catalina Sky Sur-

vey images taken with the 0.68-m Schmidt telescope (discovery observation tabulated below), the twilight images showing a coma 15′′ across. Followup 30-s exposures by E. Beshore with the Mt. Lemmon 1.5-m reflector on Oct. 10.5 UT show a bright, elongated nuclear condensation and a tail ≈ 20′′ long in p.a. ≈ 300^{o}. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, G. Sostero and E. Guido (Remanzacco, Italy, 0.45-m f/4.4 reflector) write that 70 co-added 30-s images from Oct. 10.13–10.16 show a concentrated coma with a diameter of nearly 18′′ of total mag ∼ 16.3, elongated toward the northwest, with a narrow tail extending nearly 15′′ toward p.a. 300^{o}. L. Buzzi (Varese, Italy, 0.60-m f/4.6 reflector) reports that, on his images from Oct. 11.2, the object is elongated 10′′ in p.a. 310^{o}. D. T. Durig (Sewanee, TN, U.S.A., 0.30-m f/3 Schmidt-Cassegrain reflector) notes that twenty stacked 2-min images taken on Oct. 11.4 show a 15′′–17′′ coma that is asymmetrical in p.a. 310^{o}. J. E. McGaha (Tucson, AZ, U.S.A., 0.62-m f/5.1 reflector) adds that his 120-s images from Oct. 11.48 show a 6′′ coma that is elongated toward a short 8′′ tail in p.a. 305^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Oct. 9.50984  10^{h}05^{m}54.^{s}70  +22^{o}37′06′′.5  16.9

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2007-T76.

T = 2007 Sept.18.999 TT  ω = 358.^{o}436
e = 0.78882  Ω = 4.161 }2000.0
q = 0.69221 AU  i = 9.979
a = 3.27775 AU n^{o} = 0.166089 P = 5.9 years

V458 VULPECULAE Visual magnitude estimates (cf. IAUC 8863), some forwarded by E.

Broens, Mol, Belgium, on behalf of the Werkgroep Veranderlijke Sterren, Vereniging Voor Sterrenkunde: Aug. 26.821 UT, 10.7 (K. Hornoch, Vohancice, Czech Republic); 29.843, 11.0 (P. Cloesen, Beringen, Belgium); Sept. 4.874, 11.2 (F. Van Loo, Genk, Belgium); 12.912, 11.3 (A. Diepvens, Balen, Belgium); 16.835, 11.6 (H. Hautecler, Boutersem, Belgium); 20.784, 12.0 (M. Lehky, Hradec Kralove, Czech Republic); 24.847, 11.8 (E. Muyllaert, Oostende, Belgium).

2007 October 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green